OpenStudy (anonymous):

x(d^2y/dx^2)+dy/dx=1

OpenStudy (anonymous):

\[x \frac{ d ^{2}y }{ dx ^{2} }+\frac{ dy }{ dx }=1\]

OpenStudy (anonymous):

forgot to say where \[\rho=\frac{ dy }{ dx }\]

OpenStudy (anonymous):

i am getting \[y=x+c _{1}Inx+c _{2}\] am i on the right track?

OpenStudy (anonymous):

firstly its the linear form of diff....eauation that will be used

OpenStudy (anonymous):

explain further what do u mean by that

OpenStudy (nubeer):

which approach did u use to solve the question.. i mean the method.

OpenStudy (anonymous):

i used substitution p=dy/dx and d^2x/dx^2=dp/dx

OpenStudy (anonymous):

what do u think @nubeer

OpenStudy (watchmath):

first solve \(x\rho'+\rho=1\) for \(\rho\) then solve \(dy/dx=\rho\)

OpenStudy (callisto):

@Moyo30 Actually, you got it right :\

OpenStudy (anonymous):

yep, divide through by x and get \[\frac{ d \rho }{ dx}+\frac{ \rho }{ x }=\frac{ 1 }{ x}\]

OpenStudy (anonymous):

Thx @Callisto

OpenStudy (callisto):

Welcome.. I didn't help though!!

OpenStudy (anonymous):

u answered my question.

OpenStudy (anonymous):

all i wanted to know was am i on the rite track

OpenStudy (callisto):

:)